Mike and Dan bought a foot long sub sandwich and decided that it would be fair for Dan to have a third more of the sandwich than Mike. Ho…
Found 2 solutions by stanbon, JulietG:
Answer by stanbon(75887)

Mike and Dan bought a foot long sub sandwich and decided that it would be fair for Dan to have a third more of the sandwich than Mike.
----------------------
Let Mike have "x":
Then Dan must have "(4/3)x"
------
Equation:
x + (4/3)x = 12 inches
-----
(7/3)x = 12
x = (3/7)12 = 36/7 = 5 1/7 inches (Mike's share)
(4/3)x = (4/3)(36/7) = 48/7 == 6 6/7 inches (Dan's share)
---------------
Cheers,
Stan H.
---------------

Answer by JulietG(1812)

M + D = 1
D = 1.33 M
Substitute the value of D from the second equation into the first.
M + D = 1
M + (1.33M) = 1
2.33M = 1
Divide each side by 2.33
1 / 2.33 = ~ 0.43
Dan gets 57% of the sub and Mike gets 43%
For a 12" sub, that's 6.84" for Dan and 5.16" for Mike. (Approximately due to rounding.)